# Detecting Misbehaviors of Large Vision-Language Models by Evidential Uncertainty Quantification

### v1

@@ Abstract
Large vision-language models (LVLMs) have shown substantial advances in multimodal understanding and generation. However, when presented with incompetent or adversarial inputs, they frequently produce unreliable or even harmful content, such as fact hallucinations or dangerous instructions. This misalignment with human expectations, referred to as misbehaviors of LVLMs, raises serious concerns for deployment in critical applications. These misbehaviors are found to stem from epistemic uncertainty, specifically either conflicting internal knowledge or the absence of supporting information. However, existing uncertainty quantification methods, which typically capture only overall epistemic uncertainty, have shown limited effectiveness in identifying such issues. To address this gap, we propose Evidential Uncertainty Quantification (EUQ), a fine-grained method that captures both information conflict and ignorance for effective detection of LVLM misbehaviors. In particular, we interpret features from the model output head as either supporting (positive) or opposing (negative) evidence. Leveraging Evidence Theory, we model and aggregate this evidence to quantify internal conflict and knowledge gaps within a single forward pass. We extensively evaluate our method across four categories of misbehavior, including hallucinations, jailbreaks, adversarial vulnerabilities, and out-of-distribution (OOD) failures, using state-of-the-art LVLMs, and find that EUQ consistently outperforms strong baselines, showing that hallucinations correspond to high internal conflict and OOD failures to high ignorance. Furthermore, layer-wise evidential uncertainty dynamics analysis helps interpret the evolution of internal representations from a new perspective. The source code is available at https://github.com/HT86159/EUQ.
@@ 1 Introduction
Large Vision-Language Models (LVLMs) (Liu et al., 2024c; Bai et al., 2025; Wu et al., 2024b) have demonstrated remarkable capabilities in multimodal understanding and context-aware reasoning across a variety of vision-language tasks (Ngiam et al., 2011; Chen et al., 2020). Nevertheless, their outputs can become unreliable or even harmful manner when faced with complex problem-solving scenarios or adversarial inputs. Such challenges often lead to issues such as unfaithful hallucinations (Biten et al., 2022; Li et al., 2023), security risks through jailbreaks (Qi et al., 2024; Gong et al., 2025), adversarial vulnerabilities (Fang et al., 2024; Ge et al., 2023), and failures to generalize out-of-distribution (OOD) (Yang et al., 2024; Xu et al., 2025). These misbehaviors indicate that current LVLMs are not yet fully aligned with human expectations (Herce Castañón et al., 2019; Denison et al., 2018; Qi et al., ). As a result, such failures significantly hinder their deployment in critical applications, their deployment in critical domains, such as automated financial trading (Xiong et al., 2025), autonomous driving (Grigorescu et al., 2020) and medical diagnosis (Kumar et al., 2023), remains a significant challenge. This underscores the urgent need for effective detection and mitigation methods to enhance model trustworthiness.
The connection between such misbehaviors and model uncertainty has been widely recognized (Amodei et al., 2016; ISO, 2022). Our focus mainly lies on a significant and reducible component, epistemic uncertainty, which the limitation in model knowledge captured by its parameters. This uncertainty has long been understood to originate from two primary sources (Denœux et al., 2020): the presence of conflicting information and the absence of supporting information. For instance, the top case in Figure 1 illustrates the former, the model correctly identifies both the text and the background image, yet their semantic inconsistency leads to a response that casts doubt on the input. In contrast, the bottom example shows the latter, with the model perceiving color and shape but expressing “cannot immediately identify” and resorting to “guessing” due to missing information.
@@
While misbehaviors in LVLMs often arise from internal conflicts or knowledge absence, existing uncertainty quantification (UQ) approaches focus on the total predictive uncertainty, but fail to explicitly capture such underlying causes. Most classical uncertainty quantification (UQ) methods, such as Bayesian approaches (MacKay, 1992; Blundell et al., 2015) and their variants (Gal and Ghahramani, 2016; Lakshminarayanan et al., 2017; Maddox et al., 2019), as well as internal methods of deterministic models (Sensoy et al., 2018; Malinin and Gales, 2019), are challenging to apply to LVLMs due to their substantial computational overhead. As a result, recent efforts predominantly adopt test-time sampling strategies. A typical strategy estimates overall epistemic uncertainty from token-level probabilities of a single output (Kadavath et al., 2022; Malinin and Gales, 2021), while subsequent extensions evaluate semantic variability across multiple generations (Farquhar et al., 2024; Manakul et al., 2023). Another line of work encourages models to verbalize their confidence (Xiong et al., ; Lin et al., ). However, such uncertainty is often unstable and uncalibrated, as LVLMs lack strong metacognitive capabilities, i.e., they struggle to reliably recognize and express their own uncertainty.
@@
To this end, we propose Evidential Uncertainty Quantification (EUQ), which enables effective and computationally efficient detection of model misbehaviors. To the best of our knowledge, this is the first attempt to explicitly characterize two types of epistemic uncertainty in LVLMs, conflict (CF) and ignorance (IG). CF quantifies the degree of contradiction among evidence in model predictions, while IG measures the lack of information available to the model. Specifically, we draw inspiration from the interpretation of linear projection as evidence fusion (Denœux, 2019). Evidence is then constructed from the pre-logits features of the LVLM output head, which provide high-level signals directly linked to the model’s decisions (Zhao et al., 2024), to quantify uncertainty. We then apply basic belief assignment (BBA), which distributes belief masses over hypotheses, to convert them into evidence weights. These weights are then decomposed into positive and negative components, which represent support and contradiction to the model’s decision. The refined evidence weights are fused using Dempster’s rule of combination (Shafer, 1976), yielding CF from the conflict between positive and negative evidence and IG from the missing information in the fused evidence. As shown in Figure 1, $\mathrm{CF}$ primarily highlights concrete objects (e.g., diaper), whereas $\mathrm{IG}$ captures both objects and modifiers (e.g., vaguely). Tokens corresponding to the final decision exhibit high CF and IG values, indicating that these measures effectively capture the sources of uncertainty. We evaluate our method on misbehavior detection across four scenarios, encompassing hallucinations, jailbreaks, adversarial vulnerabilities, and OOD failures. Comprehensive experiments on DeepSeek-VL2-Tiny (Wu et al., 2024b), Qwen2.5-VL-7B (Bai et al., 2025), InternVL2.5-8B (Chen et al., 2024), and MoF-Models-7B (Tong et al., 2024) show that CF and IG consistently outperform strong baselines, achieving relative improvements of 10.4%/7.5% AUROC and 5.3%/5.5% AUPR. Empirical analysis further reveals that hallucinations correspond to high internal conflict, whereas OOD failures correspond to high ignorance. Our contributions are summarized as follows:
@@
  • We identify that diverse misbehaviors in LVLMs primarily stem from two types of epistemic uncertainty: internal contradictions and missing supporting information. To address this, we propose a computationally efficient, DST-based detection method that captures these fine-grained uncertainties in a single forward pass.
@@
  • Extensive experiments on four advanced LVLMs across four behavior scenarios demonstrate that our method consistently outperforms strong baselines, yielding improvements of 10.4%/7.5% in AUROC and 5.3%/5.5% in AUPR.
@@ 2.1 Misbehaviors in LVLMs
Hallucination in LVLMs denotes mismatches between visual inputs and generated text (Liu et al., 2024b). It can be categorized into three types: object hallucination, describing nonexistent objects (Biten et al., 2022; Hu et al., ; Li et al., 2023); relation hallucination, misrepresenting spatial or semantic relations (Wu et al., 2024a); attribute hallucination, assigning wrong properties to visual entities (Liu et al., 2024a). Jailbreak refers to eliciting harmful behaviors misaligned with human intent, often triggered by visual perturbations (Carlini et al., 2023), exposing vulnerabilities beyond typical prediction errors. Such attacks are broadly categorized into optimization-based methods, which iteratively modify inputs via gradients or search strategies (Qi et al., 2024; Wang et al., 2024b; Bailey et al., 2024), and generation-based methods, which embed harmful typography on the clean images (Gong et al., 2025; Li et al., 2024; Goh et al., 2021; Shayegani et al., ). Adversarial vulnerability in vision models stems from imperceptible adversarial perturbations that induce incorrect predictions (Szegedy et al., 2014; Goodfellow et al., 2014; Fang et al., 2024; Ge et al., 2023). Recent work shows that LVLMs inherit this weakness (Sheng et al., 2021; Zhao et al., 2023; Wang et al., 2024a), remaining susceptible to visual perturbations despite their multimodal nature. OOD failure refers to the inability of a model to handle inputs outside the training distribution, challenging accurate recognition (Kim et al., 2025; Han et al., ). Prior work has focused on multimodal models, like CLIP (Radford et al., 2021), for detecting inputs outside the in-distribution (ID) (Ming et al., 2022; Jiang et al., ; Cao et al., 2024). Although OOD in LVLMs is less studied, recent work defines ID inputs as standard data and OOD inputs as style or quality shifts (Kim et al., 2025; Xu et al., 2025).
@@ 2.2 Uncertainty Quantification for LVLMs
Classical UQ methods, such as Bayesian approaches (MacKay, 1992; Blundell et al., 2015) and their variants (Gal and Ghahramani, 2016; Lakshminarayanan et al., 2017; Maddox et al., 2019), are computationally expensive and thus difficult to apply to LVLMs. Deterministic methods, such as (Malinin and Gales, 2019) and  (Sensoy et al., 2018; Li et al., 2025), the latter following an evidential framework, still require model training. In contrast, our approach performs evidence modeling and aggregation at inference, producing richer uncertainty measures without additional training, making it well suited for LVLMs. Thus, this section reviews prior work on UQ for LVLMs.
Token-wise probability-based methods estimate uncertainty within a single generation using log-likelihoods (Kadavath et al., 2022; Guerreiro et al., 2023; Duan et al., 2024) and entropy measures (Malinin and Gales, 2021). However, softmax outputs tend to be overconfident (Gal and Ghahramani, 2016; Guo et al., 2017), resulting in miscalibrated uncertainty. Sampling-based methods further estimate uncertainty by evaluating variability semantics across multiple generations. (Lin et al., 2023) estimates uncertainty via pairwise similarities and a graph Laplacian. (Farquhar et al., 2024) proposes semantic entropy to detect confabulations, utilizing external models to evaluate semantic equivalence. Other works (Raj et al., 2023; Manakul et al., 2023) design task-specific prompts and use auxiliary LLMs to assess semantic consistency. Regardless, these methods are computationally expensive due to repeated inference and heavily depend on auxiliary models. Verbal elicitation approaches, completely independent of output probabilities, estimate a model’s uncertainty by prompting it to express self-assessments in natural language. (Lin et al., ) introduces verbalization probability and demonstrates its alignment with model logits after fine-tuning. Subsequent studies (Tian et al., 2023; Zhou et al., 2023; Xiong et al., 2024) focus on prompting strategies, such as employing Chain-of-Thought (CoT) (Kojima et al., 2022) to improve verbalized uncertainty, which depends heavily on the model’s compliance with prompts (Kapoor et al., 2024).
@@ 4.1 Experimental Settings
Hallucination (Li et al., 2023) 1000 Multiple-choice
@@
Jailbreak (Li et al., 2024) 200 Open-ended
@@
Datasets We evaluate our method and baselines on hallucination scenarios using POPE (Li et al., 2023) and R-Bench (Wu et al., 2024a), focusing on object and relation hallucinations. For jailbreak scenarios, we evaluate a range of jailbreak attacks, including FigStep (Gong et al., 2025), Hades (Li et al., 2024), and VisualAdv (Qi et al., 2024). We further simulate typographic attacks following the protocol of (Goh et al., 2021). For adversarial scenarios, we employ two state-of-the-art attacks: ANDA (Fang et al., 2024) and PGN (Ge et al., 2023). For OOD failures, we use the dataset from (Xu et al., 2025) [^1] https://huggingface.co/datasets/AI-Secure/MMDecodingTrust-I2T. All tasks include a valid correct answer, and all provided options are reasonable.
@@
Baselines We compare against four baselines: two sampling-based methods: self-consistency (SC) (Wang et al., ), semantic entropy (SE) (Farquhar et al., 2024); and two probability-based methods—predictive entropy (PE) (Kadavath et al., 2022) and its length-normalized variant (LN-PE) (Malinin and Gales, 2021). Additionally, we include HiddenDetect (Jiang et al., 2025), originally designed for jailbreak detection, which relies on refusal cues and can also be applied to other misbehaviors.
@@
Evaluation Metric for Detection We use the Area Under the ROC Curve (AUROC) to evaluate detection performance, measuring the ability to rank correct (low uncertainty) above incorrect generations. The Area Under the Precision-Recall Curve (AUPR) (Davis and Goadrich, 2006) is also reported to address data imbalance from rare misbehavior cases.

### v2

@@ Abstract
Large vision-language models (LVLMs) have achieved substantial advances in multimodal understanding. However, when presented with challenging or distribution-shifted inputs, they frequently produce unreliable or even harmful content, such as hallucinations or toxic responses. We refer to such misalignments with human expectations as misbehaviors of LVLMs, which raise serious concerns for their deployment in critical applications. Existing research have disclosed that such misbehaviors are closely linked to model uncertainty. We find they primarily stem from two distinct sources of epistemic uncertainty: internal contradictions (conflict) and the absence of supporting information (ignorance). While existing uncertainty quantification methods typically capture only total predictive uncertainty, they struggle to distinguish between these underlying causes. To address this gap, we propose Evidential Uncertainty Quantification (EUQ), a training-free framework that explicitly decomposes epistemic uncertainty into conflict (CF) and ignorance (IG). Specifically, we interpret features from the model output head as either supporting (positive) or opposing (negative) evidence. Leveraging Dempster-Shafer Theory of belief functions, we aggregate this evidence to quantify internal conflict and knowledge gaps within a single forward pass. We extensively evaluate EUQ across four misbehavior categories, including hallucinations, jailbreaks, adversarial vulnerabilities, and out-of-distribution (OOD) failures using state-of-the-art LVLMs. Experimental results demonstrate that EUQ consistently outperforms strong baselines, achieving relative improvements of up to 10.5% in AUROC. Our evaluation further reveals that hallucinations correspond to high internal conflict and OOD failures to high ignorance. Furthermore, a layer-wise evidential uncertainty dynamics analysis provides a novel perspective on the evolution of internal representations. The source code is available at https://github.com/HT86159/EUQ.
@@ 1 Introduction
Large Vision-Language Models (LVLMs) (Liu et al., 2024c; Bai et al., 2025; Wu et al., 2024b) have demonstrated remarkable capabilities in multimodal understanding and context-aware reasoning across a variety of vision-language tasks (Ngiam et al., 2011; Chen et al., 2020). Nevertheless, their outputs can become unreliable or even harmful when faced with challenging, distribution-shifted, or adversarial inputs. Such challenges often lead to issues such as unfaithful hallucinations (Biten et al., 2022; Li et al., 2023b), security risks through jailbreaks (Qi et al., 2024; Gong et al., 2025), adversarial vulnerabilities (Fang et al., 2024; Ge et al., 2023), and failures to generalize out-of-distribution (OOD) (Yang et al., 2024; Xu et al., 2025). These misbehaviors indicate that current LVLMs are not yet fully aligned with human expectations (Bengio et al., 2025; Feng et al., 2026). As a result, such failures significantly hinder their deployment in critical applications, such as identity authentication (Li et al., 2023a), autonomous driving (Grigorescu et al., 2020), and medical diagnosis (Kumar et al., 2023). This underscores the urgent need for effective detection and mitigation methods to enhance model trustworthiness.
The connection between such misbehaviors and model uncertainty has been widely recognized (Farquhar et al., 2024; Liao et al., ; ISO, 2022). Our focus mainly lies on a significant and reducible component, epistemic uncertainty, which is the limitation in model knowledge captured by its parameters. This uncertainty has long been understood to originate from two primary sources (Denœux et al., 2020): the presence of conflicting information and the absence of supporting information. For instance, the top case in Figure 1 illustrates the former; the model correctly identifies both the text and the background image, yet their semantic inconsistency leads to a response that casts doubt on the input. In contrast, the bottom example shows the latter, with the model perceiving color and shape but expressing “cannot immediately identify” and resorting to “guessing” due to missing information.
@@
While misbehaviors in LVLMs often arise from internal conflicts or knowledge absence, existing uncertainty quantification (UQ) approaches focus on the total predictive uncertainty, but fail to explicitly capture such underlying causes. Most classical uncertainty quantification (UQ) methods, such as Bayesian approaches (MacKay, 1992; Blundell et al., 2015) and their variants (Gal & Ghahramani, 2016; Lakshminarayanan et al., 2017; Maddox et al., 2019), as well as internal methods of deterministic models (Sensoy et al., 2018; Malinin & Gales, 2019), are challenging to apply to LVLMs due to their substantial computational overhead. As a result, recent efforts predominantly adopt test-time sampling strategies. A typical strategy estimates overall epistemic uncertainty from token-level probabilities of a single output (Kadavath et al., 2022; Malinin & Gales, 2021), while subsequent extensions evaluate semantic variability across multiple generations (Farquhar et al., 2024; Manakul et al., 2023; Chen et al., 2025). Another line of work encourages models to verbalize their confidence (Xiong et al., ; Lin et al., ). However, such uncertainty is often unstable and uncalibrated, as LVLMs lack strong metacognitive capabilities, i.e., they struggle to reliably recognize and express their own uncertainty.
@@
To this end, we propose Evidential Uncertainty Quantification (EUQ), which enables effective and computationally efficient detection of model misbehaviors. To the best of our knowledge, this is the first attempt to explicitly characterize two types of epistemic uncertainty in LVLMs, conflict (CF) and ignorance (IG). CF quantifies the degree of contradiction among evidence in model predictions, while IG measures the lack of information available to the model. Specifically, we draw inspiration from the interpretation of linear projection as evidence fusion (Denœux, 2019; Huang et al., 2025). Evidence is then constructed from the pre-logits features of the LVLM output head, which provide high-level signals directly linked to the model’s decisions (Zhao et al., 2024), to quantify uncertainty. We then apply basic belief assignment (BBA), which distributes belief masses over hypotheses, to convert them into evidence weights. These weights are then decomposed into positive and negative components, which represent support and contradiction to the model’s decision. The refined evidence weights are fused using Dempster’s rule of combination (Shafer, 1976), yielding CF from the conflict between positive and negative evidence and IG from the missing information in the fused evidence. As shown in Figure 1, $\mathrm{CF}$ primarily highlights concrete objects (e.g., diaper), whereas $\mathrm{IG}$ captures both objects and modifiers (e.g., vaguely). Tokens corresponding to the final decision exhibit high CF and IG values, indicating that these measures effectively capture the sources of uncertainty. We evaluate our method on misbehavior detection across four scenarios, encompassing hallucinations, jailbreaks, adversarial vulnerabilities, and OOD failures. Comprehensive experiments on DeepSeek-VL2-Tiny (Wu et al., 2024b), Qwen2.5-VL-7B (Bai et al., 2025), InternVL2.5-8B (Chen et al., 2024), and MoF-Models-7B (Tong et al., 2024) show that CF and IG consistently outperform strong baselines, achieving relative improvements of 10.4%/7.5% AUROC and 5.3%/5.5% AUPR. Empirical analysis further reveals that hallucinations correspond to high internal conflict, whereas OOD failures correspond to high ignorance. Our contributions are summarized as follows:
@@
  • We identify that diverse misbehaviors in LVLMs primarily stem from two types of epistemic uncertainty: internal contradictions and missing supporting information. To address this, we propose a computationally efficient detection method based on Dempster-Shafer Theory (DST). It captures these fine-grained uncertainties in a single forward pass.
@@
  • Extensive experiments on four advanced LVLMs using our proposed Misbehavior-Bench [^1] https://huggingface.co/datasets/thuang5288/Misbehavior-Bench demonstrate that the method consistently outperforms strong baselines. It yields improvements of 10.4%/7.5% in AUROC and 5.3%/5.5% in AUPR.
@@ 2.1 Misbehaviors in LVLMs
Hallucination in LVLMs denotes mismatches between visual inputs and generated text (Liu et al., 2024b). It can be categorized into three types: object hallucination, describing nonexistent objects (Biten et al., 2022; Hu et al., ; Li et al., 2023b); relation hallucination, misrepresenting spatial or semantic relations (Wu et al., 2024a); attribute hallucination, assigning wrong properties to visual entities (Liu et al., 2024a). Jailbreak refers to eliciting harmful behaviors misaligned with human intent, often triggered by visual perturbations (Carlini et al., 2023), exposing vulnerabilities beyond typical prediction errors. Such attacks are broadly categorized into optimization-based methods, which iteratively modify inputs via gradients or search strategies (Qi et al., 2024; Wang et al., 2024b; Bailey et al., 2024), and generation-based methods, which embed harmful typography on the clean images (Gong et al., 2025; Li et al., 2024a; Goh et al., 2021; Shayegani et al., ). Adversarial vulnerability in vision models stems from imperceptible adversarial perturbations that induce incorrect predictions (Szegedy et al., 2014; Li et al., 2023a; Fang et al., 2024; Ge et al., 2023; Qin et al., 2025). Recent work shows that LVLMs inherit this weakness (Sheng et al., 2021; Zhao et al., 2023; Wang et al., 2024a), remaining susceptible to visual perturbations despite their multimodal nature. OOD failure refers to the inability of a model to handle inputs outside the training distribution, challenging accurate recognition (Kim et al., 2025; Han et al., ). Prior work has focused on multimodal models, like CLIP (Radford et al., 2021), for detecting inputs outside the in-distribution (ID) (Ming et al., 2022; Jiang et al., ; Cao et al., 2024). Although OOD in LVLMs is less studied, recent work defines ID inputs as standard data and OOD inputs as style or quality shifts (Kim et al., 2025; Xu et al., 2025).
@@ 2.2 Uncertainty Quantification for LVLMs
Classical UQ methods, such as Bayesian approaches (MacKay, 1992; Blundell et al., 2015) and their variants (Gal & Ghahramani, 2016; Lakshminarayanan et al., 2017; Maddox et al., 2019), are computationally expensive and thus difficult to apply to LVLMs. Deterministic methods, such as (Malinin & Gales, 2019) and  (Sensoy et al., 2018; Li et al., 2025), the latter following an evidential framework, still require model training. In contrast, our approach performs evidence modeling and aggregation at inference, producing richer uncertainty measures without additional training, making it well suited for LVLMs. Thus, this section reviews prior work on UQ for LVLMs.
Token-wise probability-based methods estimate uncertainty within a single generation using log-likelihoods (Kadavath et al., 2022; Guerreiro et al., 2023; Duan et al., 2024) and entropy measures (Malinin & Gales, 2021). However, softmax outputs tend to be overconfident (Gal & Ghahramani, 2016; Guo et al., 2017), resulting in miscalibrated uncertainty. Sampling-based methods further estimate uncertainty by evaluating variability semantics across multiple generations. (Lin et al., 2023) estimates uncertainty via pairwise similarities and a graph Laplacian. (Farquhar et al., 2024) proposes semantic entropy to detect confabulations, utilizing external models to evaluate semantic equivalence. Other works (Raj et al., 2023; Manakul et al., 2023) design task-specific prompts and use auxiliary LLMs to assess semantic consistency. Regardless, these methods are computationally expensive due to repeated inference and heavily depend on auxiliary models. Verbal elicitation approaches, completely independent of output probabilities, estimate a model’s uncertainty by prompting it to express self-assessments in natural language. (Lin et al., ) introduces verbalization probability and demonstrates its alignment with model logits after fine-tuning. Subsequent studies (Tian et al., 2023; Zhou et al., 2023; Xiong et al., 2024) focus on prompting strategies, such as employing Chain-of-Thought (CoT) (Kojima et al., 2022) to improve verbalized uncertainty, which depends heavily on the model’s compliance with prompts (Kapoor et al., 2024).
@@ 4.1 Experimental Settings
Hallucination (Li et al., 2023b) 1000 Multiple-choice
@@
Jailbreak (Li et al., 2024b) 200 Open-ended
@@
Datasets We evaluate our method and baselines on hallucination scenarios using POPE (Li et al., 2023b) and R-Bench (Wu et al., 2024a), focusing on object and relation hallucinations. For jailbreak scenarios, we evaluate a range of jailbreak attacks, including FigStep (Gong et al., 2025), Hades (Li et al., 2024b), and VisualAdv (Qi et al., 2024). We further simulate typographic attacks following the protocol of (Goh et al., 2021). For adversarial scenarios, we employ two state-of-the-art attacks: ANDA (Fang et al., 2024) and PGN (Ge et al., 2023). For OOD failures, we use the dataset from (Xu et al., 2025). The aforementioned scenarios and data sources constitute our Misbehavior-Bench. All tasks include a valid, correct answer, and all provided options are reasonable.
@@
Baselines We compare against four baselines: two sampling-based methods: self-consistency (SC) (Wang et al., ), semantic entropy (SE) (Farquhar et al., 2024); and two probability-based methods—predictive entropy (PE) (Kadavath et al., 2022) and its length-normalized variant (LN-PE) (Malinin & Gales, 2021). Additionally, we include HiddenDetect (Jiang et al., 2025), originally designed for jailbreak detection, which relies on refusal cues and can also be applied to other misbehaviors.
@@
Evaluation Metric for Detection We use the Area Under the ROC Curve (AUROC) to evaluate detection performance, measuring the ability to rank correct (low uncertainty) above incorrect generations. The Area Under the Precision-Recall Curve (AUPR) (Davis & Goadrich, 2006) is also reported to address data imbalance from rare misbehavior cases.